# Measuring the Dunkelflaute: How (not) to analyze variable renewable energy shortage

### v1

@@ 2.1 Drought definitions
VRE availability factors. When exploring VRE droughts, i.e., VRE shortage focusing on the energy supply side, one approach is to detect periods with low wind speeds or solar irradiation. According data can originate from measurements (in-situ or by satellites) or meteorological models. Alternatively, drought analyses can use VRE availability time series, also referred to as capacity factors, either derived from measurements or converted from modeled weather variables. These indicate how much electricity (e.g., in megawatt hours) a VRE source can generate at a given time step $t$, normalized by its installed generation capacity (e.g., in megawatt). In other words, an availability factor is a measure of a VRE generator’s capture efficiency of the meteorological resource nuno_simulation_2018 . While our analysis focuses on discrete time series of VRE availability factors $\phi_{t}^{VRE}\in[0,1]$, the discussed methods generally apply to all of the above-mentioned data types. For clarity, we start with a single technology and one region, and later turn to renewable technology portfolios and larger regions.
Drought threshold. Given a VRE availability time series, VRE droughts can generally be defined as periods below a certain drought qualification threshold $\tau\in[0,1]$ (gray lines in Figure 2).
(a) Drought windows with fixed duration below a drought threshold $\tau=0.1$.
(b) Drought events with variable duration below two drought thresholds $\tau_{1}=0.1$ and $\tau_{1}=0.2$. Using higher thresholds generally leads to longer, but fewer, events.
@@
Drought definition: window perspective. There are different definitions for VRE drought periods. One option is searching for ’drought windows’ with predefined duration raynaud_energy_2018 ; kapica_potential_2024 ; rinaldi_wind_2021  [^1] These studies search for periods of consecutive daily averages below some threshold with a cut-off every 24 hours., e.g., a given number of time steps below a given threshold (Figure 2a). Implementing a respective search algorithm for a specified window duration is straightforward. Yet, droughts identified via predefined windows may be shorter than actual droughts (Figure 2a). Further, consecutive windows that form one longer combined drought may be counted individually (Figure 2a, right windows). Thus, this definition systematically underestimates drought duration and overestimates the number of droughts.
Drought definition: event perspective. An alternative approach is searching for complete ’drought events’ below a threshold, which last as long as a drought qualification criterion is met. That is, drought duration is not predefined by setting a window length, but, for example, the result of a search on how many consecutive time steps are below a given threshold yevyevich_objective_1967 ; fleig_global_2006 ; potisomporn_extreme_2024 . This ensures that identified events are completely captured, and consecutive droughts are properly accounted for (Figure 2b). We thus consider the drought event perspective to be a superior concept for identifying VRE droughts, and focus on it henceforward.
Threshold intuition. Higher thresholds generally identify less severe, but longer-lasting drought events. This tends to reduce the number of obtained droughts, as events that are considered to be separate for lower thresholds may integrate into combined ones (Figure 2b).
@@ 2.2 Drought identification methods
Identification method: CBT. Numerous methods exist for identifying VRE droughts. We focus on a selection [^2] We limit our selection to methods that appear useful for investigating VRE droughts. Other methods in different bodies of literature can be found, e.g, the Inter-event time method that was developed for hydropower zelenhasic_method_1987 and was recently evaluated in potisomporn_extreme_2024 . in the following (Figure 3). A Constantly-Below-Threshold (CBT) event relates to a period of consecutive time steps during which availability factors are constantly below a threshold $\tau$ guerrero-salazar_analysis_1975 ; leahy_persistence_2013 ; cannon_using_2015 ; patlakas_low_2017 ; ohlendorf_frequency_2020 ; potisomporn_spatial_2022 ; potisomporn_extreme_2024 ; mayer_probabilistic_2023 . In the hydrologic literature, this method has been first introduced as ’runs analysis’ yevyevich_objective_1967 . The boolean function $\psi_{t}^{CBT}$ determines whether time step $t$ qualifies as drought (1) or not (0):
$\psi_{t}^{CBT}=\begin{cases}1&\text{if $\phi_{t}^{VRE}<\tau\ \ \forall\ \ t$}\\ 0&\text{if $\phi_{t}^{VRE}\geq\tau\ \ \forall\ \ t$}\end{cases}$ (1)
@@
Figure 3: The upper panels shows exemplary VRE availability time series (bold line) and identification methods CBT, FMBT, VMBT, and SPA. The gray line represents the drought threshold $\tau=0.1$. Green and orange areas highlight the energy deficit obtained by CBT and FMBT, respectively. The lower panel’s horizontal bars illustrate the varying event duration $\delta$ identified by each method.
@@
CBT duration & energy deficit. The duration of a CBT event $\delta^{CBT}$ is the number of consecutive time steps with $\psi_{t}^{CBT}=1$. Suppose $t_{k}$ and $t_{l}$ are the first and the last time steps of an event. Then the event’s cumulative energy, also referred to as energy deficit or drought severity yevyevich_objective_1967 ; potisomporn_extreme_2024 , is defined as follows:
$\epsilon^{CBT}=\sum_{t_{k}}^{t_{l}}\tau-\phi_{t}^{VRE}\ \ \forall\ \ t_{k}\leq t\leq t_{l}$ (2)
@@
Limitation CBT. The CBT notion identifies VRE droughts regardless of availability factors in adjacent periods, even if their availability is briefly above the drought threshold (Figure 3).
@@
Identification method: MBT. Addressing this limitation, Mean-Below-Threshold (MBT) events define consecutive time steps during which a moving average of availability factors $\Phi_{t}^{VRE}(\omega)$ is below a threshold raynaud_energy_2018 ; ohlendorf_frequency_2020 ; kaspar_climatological_2019 ; mockert_meteorological_2023 ; potisomporn_extreme_2024 . This approach smooths out short-term fluctuations over the lagging averaging interval $[t-\omega,t]$ [^3] We propose using a lagging averaging interval, which is is common in time series analyses. Using a leading interval would also be possible, but appears non-intuitive. Likewise, a centered interval is also not recommended as dealing with border time steps for an uneven number of time steps is ambiguous. with a predefined duration $\omega$. Equivalent to $\psi_{t}^{CBT}$, the function $\psi_{t}^{MBT}$ evaluates each time step $t$ as follows:
@@
$\psi_{t}^{MBT}=\begin{cases}1&\text{if $\Phi_{t}^{VRE}(\omega)<\tau\ \ \forall\ \ t$}\\ 0&\text{if $\Phi_{t}^{VRE}(\omega)\geq\tau\ \ \forall\ \ t$}\end{cases}$ (3)
@@
Fixed-duration MBT. The MBT method can be applied in two different ways. First, the fixed-duration MBT (FMBT) searches for MBT events based on a fixed averaging interval $\omega^{FMBT}$ kaspar_climatological_2019 ; mockert_meteorological_2023 ; potisomporn_extreme_2024 . The duration of an FMBT event $\delta^{FMBT}$ corresponds to the number of consecutive time steps with $\psi_{t}^{FMBT}=1$ (e.g., $\delta^{FMBT}=15$ for $\omega=10$ in Figure 3). The energy deficit $\epsilon^{FMBT}$ is computed analogously to $\epsilon^{CBT}$ (i.e., using $\Phi_{t}^{VRE}(\omega)$ instead of $\phi_{t}^{VRE}$) and decreases due to the smoothing effect of the moving average. One option to avoid this is evaluating the energy deficit based on the original time series $\phi_{t}^{VRE}$ instead of $\Phi_{t}^{VRE}(\omega)$, i.e., reporting $\epsilon^{CBT}$ instead of $\epsilon^{FMBT}$ corresponding to the identified FMBT event potisomporn_extreme_2024 .
@@
Comparison CBT & FMBT. Unlike CBT, the FMBT approach allows for instances above the drought threshold before, between, and after potentially multiple CBT drought events that are filtered out by pooling them together (compare CBT and FMBT events in Figure 3) potisomporn_extreme_2024 . In the hydrology literature, such CBT events are referred to as ’mutually dependent’ fleig_global_2006 . For prolonging averaging intervals, the moving average smoothing becomes more pronounced. Hence, the choice of $\omega^{FMBT}$ determines to which extent shorter CBT events are filtered out. [^4] The FMBT method converges to CBT as the length of the averaging interval $\omega^{FMBT}$ approaches the granularity of the original time series. By design, for a given threshold, the CBT approach identifies more drought events than FMBT, whereas FMBT finds fewer but longer-lasting events. Thus, the CBT method may overestimate the number of drought events while underestimating the duration and energy deficit, which is particularly relevant for characterizing the cumulative energy of extreme VRE droughts. In contrast, the FMBT method acknowledges natural fluctuations in VRE availability by accounting for brief periods of higher availability within an overall low-availability period. FMBT therefore provides a more realistic view of prolonged VRE droughts and the corresponding power sector flexibility needs.
@@
Limitation FMBT. However, there is no clear relation between a power sector attribute and $\omega^{FMBT}$, whose parameterization remains arbitrary. VRE availability time series could be evaluated for a range of different $\omega^{FMBT}$. Yet, drought events based on different $\omega^{FMBT}$ may partially overlap, creating ambiguity in their characteristics like duration, energy deficit, and frequency. In fact, results may substantially change with varying $\omega^{FMBT}$ (compare the moving averages $\Phi_{t}^{VRE}(\omega=10)$ and $\Phi_{t}^{VRE}(\omega=50)$ in Figure 3, with the latter obtaining no drought at all). Consequently, VRE droughts identified via the FMBT approach are valid only for a specific parameterization of $\omega^{FMBT}$, limiting the applicability of this concept.
@@
Variable-duration MBT. Addressing this limitation, we introduce the variable-duration MBT (VMBT) method, which to the best of our knowledge is a novelty. It obtains unique VRE droughts, ranging from very long-lasting events, e.g., months, to such that last only a few time steps. This is achieved by deploying the MBT perspective for a large number of different averaging intervals, iteratively adjusting $\omega^{VMBT}$ in descending order. Importantly, to ensure that no long-duration events with $\delta^{VMBT}>\omega^{VMBT}$ are missed, the first iteration should start with a very high $\omega^{VMBT}$, so that the moving average $\Phi_{t}^{VRE}(\omega)>\tau$ across the entire time series (in the example shown in Figure 3, compare the moving average $\Phi_{t}^{VRE}(\omega^{VMBT}=50)$). In subsequent iterations, $\omega^{VMBT}$ is reduced and $\Phi_{t}^{VMBT}$ updated, until the first drought event is identified at $\Phi_{t}^{VMBT}<\tau$ (compare the moving average $\Phi_{t}^{VRE}(\omega=25)$ in Figure 3). Its duration equals the averaging interval $\delta^{VMBT}=\omega^{VMBT}$. The respective period is then excluded in later iterations, where $\omega^{VMBT}$ decreases further and additional (shorter) drought events are identified, and so on. For a given drought threshold $\tau$, this iterative process effectively captures all unique drought events, without any overlap or double counting, each accurately recorded with its maximum possible duration. Compared to FMBT, this renders the VMBT approach a more accurate and informative method for drought identification.
@@
Limitation VMBT. Note that the VMBT energy deficit $\epsilon^{VMBT}$ is insignificant due to the close proximity of the moving average to the threshold. Further, this method involves a trade-off between accuracy and computational efficiency. Iteratively reducing the interval of the moving average ($\omega^{VMBT}$), with an incremental change that equals the resolution of the underlying time series, ensures captures all unique drought events, but requires numerous iterations and may lead to long processing time.
@@
Identification method: SPA. Finally, the Sequent Peak Algorithm (SPA) identification method searches for the maximum cumulative energy deficit of a drought event, while accommodating intermediate periods where renewable availability may surpass the threshold. Originating in hydrology vogel_generalized_1987 ; fleig_global_2006 , SPA has been extended to other bodies of literature, such as wind energy analysis potisomporn_extreme_2024 . A similar concept has also been applied to characterizing renewable surplus events schill_residual_2014 . The cumulative energy deficit is defined as follows:
$\epsilon_{t}^{SPA}=\begin{cases}\epsilon^{SPA}_{t-1}+\tau-\phi_{t}^{VRE}&\text{if $\epsilon^{SPA}_{t-1}+\tau-\phi_{t}^{VRE}>0\ \ \forall\ \ t$}\\ 0&\text{if $\epsilon^{SPA}_{t-1}+\tau-\phi_{t}^{VRE}\leq 0\ \ \forall\ \ t$}\end{cases}$ (4)
@@
SPA duration. The SPA measures event duration $\delta^{SPA}$ from the first instance when the cumulative energy deficit of a single SPA event becomes positive to its global [^5] Here, “global” refers to the maximum of a single SPA event, not to the maximum of all potential SPA events of the underlying VRE availability time series. maximum. This maximum occurs when the VRE availability curve intersects with the threshold for the last time during an SPC event, i.e., when $\phi_{t}^{VRE}\geq\tau$ (Figure 3). This algorithm requires no iterations or assumptions on averaging intervals, and only a single straightforward computation of cumulative energy deficits for the underlying VRE availability time series. It allows for long-duration events that include brief periods with above-threshold availability.
@@
Comparison SPA & VMBT: Unlike VMBT, SPA events start when VRE availability falls below the threshold and thus depends on threshold parameterization. In contrast, the VMBT method allows for high or low availability periods before SPA events, making it less reliant on specific threshold parameterization. For a given threshold, both VMBT and SPA obtain a unique and unambiguous set of VRE drought events, and thus appear to be superior to the CBT of FMBT approaches. SPA is computationally less demanding than VMBT, as it requires no iterations over averaging intervals. Yet, SPA may not recognize minor events that follow a longer-lasting drought, which has been criticized in the literature potisomporn_extreme_2024 .
@@
Comparison duration CBT, FMBT, VMBT, SPA. As indicated by Figure 3, for a given threshold, the VMBT approach generally leads to the longest drought duration of all four identification strategies explained above. The SPA method identifies events with extended duration, albeit shorter than VMBT, followed by FMBT events with moderate duration. On the other end of the spectrum, the CBT method leads to the shortest drought duration.
@@
Limitation of energy deficits in VRE drought analysis. While often reported in other bodies of literature, energy deficits of VRE droughts identified purely based on VRE availability are hard to interpret as they do not allow for drawing direct energy system conclusions.
@@
Threshold parameterization. Note that an effective threshold parameterization is methodologically challenging and subject to user discretion (see Section 2.3 for further discussion).
2.3 Comparability and the choice of threshold

### v2

@@ 2.1 Drought definitions
When exploring VRE droughts, i.e., shortage on the energy supply side, one approach is to detect periods with low wind speeds or solar irradiation. According data can originate from measurements (in-situ or by satellites) or meteorological models. Alternatively, drought analyses can use VRE availability time series, also referred to as capacity factors, either derived from measurements or converted from modeled weather variables. These indicate how much electricity (e.g., in megawatt hours) a renewable source can generate at a given time step $t$, normalized by its installed generation capacity (e.g., in megawatt). In other words, an availability factor is a measure of a VRE generator’s capture efficiency of the meteorological resource [51]. While our analysis focuses on discrete time series of VRE availability factors $avail_{t}\in[0,1]$, the discussed methods generally apply to all of the above-mentioned data types. For clarity, we start with a single technology and one region, and later turn to renewable technology portfolios and larger regions.
Given a VRE availability time series, variable renewable energy droughts can generally be defined as periods below a certain drought qualification threshold $thres\in[0,1]$ (gray lines in Figure 2). There are different definitions for VRE drought periods. One option is searching for ’drought windows’ with a predefined duration [19, 35, 18] [^1] These studies search for periods of consecutive daily averages below some threshold with a cut-off every 24 hours., e.g., a given number of time steps below a given threshold (Figure 2a). Implementing a respective search algorithm for a specified window duration is straightforward. Yet, droughts identified via predefined windows may be shorter than actual droughts (Figure 2a). Further, consecutive windows that form one longer combined drought may be counted individually (Figure 2a, right windows). Thus, this definition systematically underestimates drought duration and overestimates the number of droughts.
(a) Drought windows with fixed duration below a drought threshold $thres=0.1$.
(b) Drought events with variable duration below two drought thresholds $thres_{1}=0.1$ and $thres_{1}=0.2$. Using higher thresholds generally leads to longer, but fewer, events.
@@
An alternative approach is searching for complete ’drought events’ below a threshold, which last as long as a drought qualification criterion is met. That is, drought duration is not predefined by setting a window length, but, for example, the result of a search on how many consecutive time steps are below a given threshold [27, 28, 17]. This ensures that identified events are completely captured, and consecutive droughts are properly accounted for (Figure 2b). We thus consider the drought event perspective to be a superior concept for identifying VRE droughts, and focus on it henceforward.
Higher thresholds generally identify less severe, but longer-lasting drought events. This tends to reduce the number of obtained droughts, as events that are considered to be separate for lower thresholds may integrate into combined ones (Figure 2b).
@@ 2.2 Drought identification methods
Numerous methods exist for identifying VRE droughts. We focus on a selection [^2] We limit our selection to methods that appear useful for investigating VRE droughts. Other methods in different bodies of literature can be found, e.g, the Inter-event time method that was developed for hydropower [52] and was recently evaluated in [17]. in the following (Figure 3). A Constantly-Below-Threshold (CBT) event relates to a period of consecutive time steps during which availability factors are constantly below a threshold $thres$ [53, 30, 24, 31, 32, 25, 17, 40]. In the hydrologic literature, this method has been first introduced as ’runs analysis’ [27]. The boolean function $VRED_{t}^{CBT}$ determines whether time step $t$ qualifies as a variable renewable energy drought (1) or not (0):
$VRED_{t}^{CBT}=\begin{cases}1&\text{if $avail_{t}<thres\ \ \forall\ \ t$}\\ 0&\text{if $avail_{t}\geq thres\ \ \forall\ \ t$}\end{cases}$ (1)
@@
Figure 3: The upper panel shows exemplary availability time series (bold line) and identification methods Constantly-Below-Threshold (CBT), FMBT, VMBT, and SPA. The gray line represents the drought threshold $thres=0.1$. Green and orange areas highlight the energy deficit obtained by CBT and FMBT, respectively. The lower panel’s horizontal bars illustrate the varying event duration $dur$ identified by each method.
@@
The duration of a CBT event $dur^{CBT}$ is the number of consecutive time steps with $VRED_{t}^{CBT}=1$. Suppose $t_{k}$ and $t_{l}$ are the first and the last time steps of an event. Then the event’s cumulative energy, also referred to as energy deficit or drought severity [27, 17], is defined as follows:
$ED_{t}^{CBT}=\sum_{t_{k}}^{t_{l}}thres-avail_{t}\ \ \forall\ \ t_{k}\leq t\leq t_{l}$ (2)
@@
The CBT notion identifies VRE droughts regardless of availability factors in adjacent periods, even if their availability is briefly above the drought threshold (Figure 3).
@@
Addressing this limitation, Mean-Below-Threshold (MBT) events define consecutive time steps during which a moving average of availability factors $MA_{t}^{avail}(intdur)$ is below a threshold [19, 32, 38, 39, 17]. This approach smooths out short-term fluctuations over the lagging averaging interval $[t-intdur,t]$ [^3] We propose using a lagging averaging interval, which is is common in time series analyses. Using a leading interval would also be possible, but appears non-intuitive. Likewise, a centered interval is also not recommended as dealing with border time steps for an uneven number of time steps is ambiguous. with a predefined duration $intdur$. Equivalent to $VRED_{t}^{CBT}$, the function $VRED_{t}^{MBT}$ evaluates each time step $t$ as follows:
@@
$VRED_{t}^{MBT}=\begin{cases}1&\text{if $MA_{t}^{avail}(intdur)<thres\ \ \forall\ \ t$}\\ 0&\text{if $MA_{t}^{avail}(intdur)\geq thres\ \ \forall\ \ t$}\end{cases}$ (3)
@@
The MBT method can be applied in two different ways. First, the fixed-duration Mean-Below-Threshold (FMBT) searches for MBT events based on a fixed averaging interval duration $intdur^{FMBT}$ [38, 39, 17]. The duration of an FMBT event $dur^{FMBT}$ corresponds to the number of consecutive time steps with $VRED_{t}^{FMBT}=1$ (e.g., $dur^{FMBT}=15$ for $intdur=10$ in Figure 3). The energy deficit $ED_{t}^{FMBT}$ is computed analogously to $ED_{t}^{CBT}$ (i.e., using $MA_{t}^{avail}(intdur)$ instead of $avail_{t}$) and decreases due to the smoothing effect of the moving average. One option to avoid this is evaluating the energy deficit based on the original time series $avail_{t}$ instead of $MA_{t}^{avail}(intdur)$, i.e., reporting $ED_{t}^{CBT}$ instead of $ED_{t}^{FMBT}$ corresponding to the identified FMBT event [17].
@@
Unlike CBT, the FMBT approach allows for instances above the drought threshold before, between, and after potentially multiple CBT drought events that are filtered out by pooling them together (compare CBT and FMBT events in Figure 3) [17]. In the hydrology literature, such CBT events are referred to as ’mutually dependent’ [28]. For prolonging averaging intervals, the moving average smoothing becomes more pronounced. Hence, the choice of $intdur^{FMBT}$ determines to which extent shorter CBT events are filtered out. [^4] The FMBT method converges to CBT as the averaging interval duration $intdur^{FMBT}$ approaches the granularity of the original time series. By design, for a given threshold, the CBT approach identifies more drought events than FMBT, whereas FMBT finds fewer but longer-lasting events. Thus, the CBT method may overestimate the number of drought events while underestimating the duration and energy deficit, which is particularly relevant for characterizing the cumulative energy of extreme VRE droughts. In contrast, the FMBT method acknowledges natural fluctuations in VRE availability by accounting for brief periods of higher availability within an overall low-availability period. FMBT therefore provides a more realistic view of prolonged VRE droughts and the corresponding power sector flexibility needs.
@@
However, there is no clear relation between a power sector attribute and $intdur^{FMBT}$, whose parameterization remains arbitrary. VRE availability time series could be evaluated for a range of different $intdur^{FMBT}$. Yet, drought events based on different $intdur^{FMBT}$ may partially overlap, creating ambiguity in their characteristics like duration, energy deficit, and frequency. In fact, results may substantially change with varying $intdur^{FMBT}$ (compare the moving averages $MA_{t}^{avail}(intdur=10)$ and $MA_{t}^{avail}(intdur=50)$ in Figure 3, with the latter obtaining no drought at all). Consequently, VRE droughts identified via the FMBT approach are valid only for a specific parameterization of $intdur^{FMBT}$, limiting the applicability of this concept.
@@
Addressing this limitation, we introduce the variable-duration Mean-Below-Threshold (VMBT) method, which to the best of our knowledge is a novelty. It obtains unique VRE droughts, ranging from very long-lasting events, e.g., months, to such that last only a few time steps. This is achieved by deploying the MBT perspective for a large number of different averaging intervals, iteratively adjusting $intdur^{VMBT}$ in descending order. Importantly, to ensure that no long-duration events with $dur^{VMBT}>intdur^{VMBT}$ are missed, the first iteration should start with a very high $intdur^{VMBT}$, so that the moving average $MA_{t}^{avail}(intdur^{VMBT})>thres$ across the entire time series (in the example shown in Figure 3, compare the moving average $MA_{t}^{avail}(intdur=50)$). In subsequent iterations, $intdur^{VMBT}$ is reduced and $MA_{t}^{avail}$ updated, until the first drought event is identified at $MA_{t}^{avail}(intdur^{VMBT})<thres$ (compare the moving average $MA_{t}^{avail}(intdur=25)$ in Figure 3). Its duration equals the averaging interval duration $dur^{VMBT}=intdur^{VMBT}$. The respective period is then excluded in later iterations, where $intdur^{VMBT}$ decreases further and additional (shorter) drought events are identified, and so on. For a given drought threshold $thres$, this iterative process effectively captures all unique drought events, without any overlap or double counting, each accurately recorded with its maximum possible duration. Compared to FMBT, this renders the VMBT approach a more accurate and informative method for drought identification.
@@
Note that the VMBT energy deficit $ED_{t}^{VMBT}$ is insignificant due to the close proximity of the moving average to the threshold. Further, this method involves a trade-off between accuracy and computational efficiency. Iteratively reducing the interval duration of the moving average ($intdur^{VMBT}$), with an incremental change that equals the resolution of the underlying time series, ensures captures all unique drought events, but requires numerous iterations and may lead to long processing time.
@@
Finally, the Sequent Peak Algorithm (SPA) identification method searches for the maximum cumulative energy deficit of a drought event, while accommodating intermediate periods where renewable availability may surpass the threshold. Originating in hydrology [54, 28], SPA has been extended to other bodies of literature, such as wind energy analysis [17]. A similar concept has also been applied to characterizing renewable surplus events [55]. The cumulative energy deficit is defined as follows:
$ED_{t}^{SPA}=\begin{cases}ED^{SPA}_{t-1}+thres-avail_{t}&\text{if $ED^{SPA}_{t-1}+thres-avail_{t}>0\ \ \forall\ \ t$}\\ 0&\text{if $ED^{SPA}_{t-1}+thres-avail_{t}\leq 0\ \ \forall\ \ t$}\end{cases}$ (4)
@@
The SPA measures event duration $dur^{SPA}$ from the first instance when the cumulative energy deficit of a single SPA event becomes positive to its global [^5] Here, “global” refers to the maximum of a single SPA event, not to the maximum of all potential SPA events of the underlying VRE availability time series. maximum. This maximum occurs when the VRE availability curve intersects with the threshold for the last time during an SPA event, i.e., when $avail_{t}\geq thres$ (Figure 3). This algorithm requires no iterations or assumptions on averaging intervals, and only a single straightforward computation of cumulative energy deficits for the underlying VRE availability time series. It allows for long-duration events that include brief periods with above-threshold availability.
@@
Unlike VMBT, SPA events start when VRE availability falls below the threshold and thus depends on threshold parameterization. In contrast, the VMBT method allows for high or low availability periods before SPA events, making it less reliant on specific threshold parameterization. For a given threshold, both VMBT and SPA obtain a unique and unambiguous set of VRE drought events, and thus appear to be superior to the CBT of FMBT approaches. SPA is computationally less demanding than VMBT, as it requires no iterations over averaging intervals. Yet, SPA may not recognize minor events that follow a longer-lasting drought, which has been criticized in the literature [17].
@@
As indicated by Figure 3, for a given threshold, the VMBT approach generally leads to the longest drought duration of all four identification strategies explained above. The SPA method identifies events with extended duration, albeit shorter than VMBT, followed by FMBT events with moderate duration. On the other end of the spectrum, the CBT method leads to the shortest drought duration.
@@
While often reported in other bodies of literature, energy deficits of VRE droughts identified purely based on VRE availability are hard to interpret as they do not allow for drawing direct energy system conclusions.
@@
Note that an effective threshold parameterization is methodologically challenging and subject to user discretion (see Section 2.3 for further discussion).
2.3 Comparability requirements and the choice of threshold